OpenStudy (anonymous):

How do I find the 5th and 6th term of this sequence: 15, 14, 12, 9, 5,...? It is neither arithmetic or geometric sequence.

OpenStudy (anonymous):

It looks like it goes by -1, -2, -3 ...

OpenStudy (anonymous):

6th term is 0

OpenStudy (anonymous):

i think the 5th term is 0.

OpenStudy (anonymous):

millie gave us 5th term

OpenStudy (anonymous):

just a slip...

OpenStudy (anonymous):

ohh, how did you get 0?

OpenStudy (anonymous):

it goes -1 -2 -3 -4 -5 next term = 5 - 5 = 0

OpenStudy (anonymous):

15 is first term 15 - 1 = 14 14-2 = 12 12-3 n = 9 etc..

OpenStudy (anonymous):

12 - 3 = 9 9 - 4 = 5

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Oh yeah. Sorry!

OpenStudy (anonymous):

okay! :D